# DP-DyLoRA: Fine-Tuning Transformer-Based Models On-Device under Differentially Private Federated Learning using Dynamic Low-Rank Adaptation

### v3

@@ IV-E DyLoRA
$\begin{split}B_{b}&=B[:,1:b]\\ A_{b}&=A[1:b,:],\end{split}$ (4)
@@
IV-F DP-DyLoRA
Here, we propose to apply DyLoRA in a federated setting with differential privacy (DP). This runs up against a problem: the choice of rank $b$ would naturally be made per client, but this would not work with DP. If different clients were to send up different-length vectors, this would immediate break DP. If instead clients padded the statistics they sent up with zeros, this would decrease the signal-to-noise-ratio on the highest ranks significantly.
DP-DyLoRA freezes all pre-trained weights and adds new trainable LoRA modules to the model and to make fine-tuning large pre-trained models more parameter-efficient. At client side, users train on their local data with a modified forward pass:
$h=W_{\text{pt}}x+\Delta Wx=W_{\text{pt}}x+B_{b}A_{b}x.$ (5)
Here, $W_{\text{pt}}$ is frozen and only $B_{b}$ and $A_{b}$ are trainable. The outputs of $W_{\text{pt}}x$ and $B_{b}A_{b}x$ are summed coordinate-wise and are given the same input for the forward pass. The updates to the trainable parameters $B_{b}$ and $A_{b}$ are then clipped and sent back to the server for aggregation and noise addition as in DP-FedAvg. The communication cost apart from the initial transfer of $W_{\text{pt}}$ is therefore equivalent to DP-LoRA with $r=\frac{r_{min}+r_{max}}{2}$ on average which is approximately half of that of DP-LoRA with $r=r_{max}$ assuming that $r_{min}=1$. Since the magnitude of the added noise grows with the number of parameters updated [16, 17, 18], DP-DyLoRA also achieves higher signal-to-noise ratio than DP-LoRA under the same DP-FL setting. Meanwhile, the same level of model expressiveness is preserved as the model architecture and number of trainable parameters remain the same for the global model.
@@ IV-F DP-DyLoRA
V Experimental Setup
@@ V-B Models
The same model is used for tasks of the same domain. Therefore, we use BERT-small [4] for experiments on Sentiment140 and R8, ViT-small [63] for CIFAR-10 and WikiArt, and DistilHuBERT [64], [65] for Speech Commands and MINDS-14. Despite the fact that these models are extremely small compared to state-of-the-art large language models with rapidly increasing sizes such as LLaMA [66], [67] with 70 billion parameters and GPT-4 [68] with 1.7 trillion parameters, models with over 20 million parameters are considered large for either on-device deployment or differentially private federated learning.
@@ V-D Non-IID Partitioning
We therefore choose to utilise Dirichlet distribution to artificially achieve non-IID label distribution for the remaining four datasets, including R8, CIFAR-10, WikiArt and MINDS-14. For these datasets, we partition by drawing from a Dirichlet distribution with $\alpha=0.1$ by default following [69, 70, 71, 72] which results in each client holding samples from very few classes.
V-E Noise Mechanism
@@
V-F Large Cohort Noise-Level Simulation
Following [45] and [75], we simulate the noise-level of larger cohort sizes with smaller ones. In practice, differentially private federated learning (DP-FL) is applied to systems with millions of clients [22]. However, it is infeasible to simulate this many clients due to resource constraints. Since a larger cohort size $C$ leads to less noise added for the same privacy guarantee, we simulate realistically large cohort size $C_{\text{large}}$ with smaller cohort size $C_{\text{small}}$. This makes our results more meaningful in terms of practical deployment of DP-FL. We follow the approach in [45] and [75] for achieving this. The noise $\sigma$ we use for simulation is then computed as $\sigma=\frac{C_{\text{small}}}{C_{\text{large}}}z_{\text{large}}\cdot S$ where $z_{\text{large}}$ is the noise multiplier calculated based on $C_{\text{large}}$.
VI Experiments
@@
VI-A Full Fine-tuning
DP-FL tends to degrade model performance due to the noise added to model updates. Previous works have proven that there is a proportional relationship between the number of updated parameters and the magnitude of the added noise [16, 17, 18]. Hence, it is particularly challenging to train large models with differential privacy. Together with the potential non-independent and identically distributed (non-IID) data distribution challenge, large transformer models are likely to fail when learning under DP-FL.
@@ VI-B Data Heterogeneity
As shown in Figure 5 and 4, we train the model on each task with both IID and non-IID data partitioning using a single combination of client number and privacy budget taken from Q1. In Figure 4, we show results with $\alpha=[0.01,0.1,1000]$ for R8, CIFAR-10, WikiArt and MINDS-14 following [69, 70, 71, 72] which are non-IID by drawing labels from a Dirichlet distribution. When $\alpha$ is set to $0.1$, there is hardly any difference from IID distribution with $\alpha=1000$ except for WikiArt on which the accuracy drops by approximately 3%. After further increasing the level of data heterogeneity by decreasing $\alpha$ to $0.01$, results are roughly the same on R8 and WikiArt. However, on CIFAR-10 and MINDS-14, model performance degrades by approximately 14% in accuracy and 10% in word error rate, respectively. This is caused by severe client drift due to data heterogeneity as shown previously in Appendix A.
For Sentiment140 and Speech Commands which are both non-IID by natural factors, we can see from Figure 5 that the model achieves similar performance on the latter with both IID and non-IID distributions. This is likely due to the similar sample distributions by class shown in Appendix A. However, on the Sentiment140 dataset, the model performs noticeably better under IID data partitioning with approximately 4% improvement. Since Sentiment140 is a binary classification dataset, some clients may only hold samples of a single class even if it is not partitioned to have a non-IID label distribution. This leads to a relatively high level of data heterogeneity which is realistic in practice due to users having different interests and habits.
@@
VI-C Parameter-efficient Fine-tuning
Recent works [30, 19, 20] have started utilising parameter-efficient fine-tuning (PEFT) methods to fine-tune transformer models with differential privacy via both central and federate learning. Apart from the obvious benefits of lower computation and communication cost, the primary motivation originates from the proven fact that fewer trainable parameters lead to better privacy-utility trade-offs [16, 17, 18].
We thereby start with comparing model training via full fine-tuning and parameter-efficient fine-tuning with the same number of clients and privacy budget. Here, we use LoRA as an example of PEFT methods as it’s empirically proven to be superior to other popular PEFT methods on natural language understanding tasks in [30] and is used in [19] for research on DP-FL.
@@ VI-C Parameter-efficient Fine-tuning
Next, we benchmark exising PEFT methods under DP-FL. We experiment with PEFT methods including Adapter, Compacter, BitFit and LoRA which are often considered in existing works on parameter-efficient fine-tuning [29, 30, 31]. Our benchmark covers three different domains including natural language understanding (NLU), computer vision (CV) and speech, similar to our previous experiments. Regarding datasets, we again choose to use Sentiment140/R8 for NLU, CIFAR-10/WikiArt for CV and Speech Commands V1/MINDS-14 for the speech domain.
Hyperparameter: For federated learning and privacy parameters, we use 1 million users, a subsampling rate of 1%, $\epsilon$=2, and $\delta$=1e-6 for all five datasets. For clipping threshold, we search over three values {0.1, 1.0, 10.0}. For Sentiment140, we train for 300 rounds and search over five learning rates {5e-2, 1e-1, 2e-1, 5e-1, 1e-0}. For R8, we train for 200 rounds and search over four learning rates {1e-1, 2e-1, 5e-1, 1e-0}. For CIFAR-10, we train for 100 rounds and search over four learning rates {2e-2, 5e-2, 1e-1, 2e-1}. For WikiArt, we train for 200 rounds and search over four learning rates {5e-2, 1e-1, 2e-1, 5e-1}. For Speech Commands V1, we train for 300 rounds and search over four learning rates {2e-1, 5e-1, 1e-0, 2e-0}. For MINDS-14, we train for 2000 rounds and search over four learning rates {2e-2, 5e-2, 1e-1, 2e-1}. Regarding parameters for DP-Adapter, DP-Compacter and DP-LoRA, we choose to use $r$=16 for all three methods and additionally $n$=8 for DP-Compacter which are derived from previous works including [29], [30].
@@
Results: Our benchmarking results covering DP-Adapter, DP-Compacter, DP-BitFit and DP-LoRA across three different domains are shown in Table IV and V. On Sentiment140, DP-Adapter achieves the best accuracy of 70.7% which is marginally higher than an accuracy of 70.4% achieved by DP-LoRA. On R8, DP-LoRA gives the best accuracy of 90.0%. The accuracy achieved by DP-Compacter and DP-BitFit are noticeably worse on the two text classification datasets. For image classification on CIFAR-10 and WikiArt, DP-LoRA achieves the best accuracy of 90.6% and 61.7% respectively, followed by 89.9% on CIFAR-10 and 49.7% on WikiArt produced by DP-BitFit. On CIFAR-10, both DP-Adapter and DP-Compacter suffer from slow and unstable convergence which subsequently leads to much worse accuracy achieved. On Speech Commands V1, DP-LoRA once again achieves the best accuracy of 92.5%. The performance of the other three DP-PEFT methods are all worse than DP-LoRA by a noticeable gap on this dataset. The word error rates (WERs) of 68.7% and 67.6% achieved by DP-Adapter and DP-Compacter respectively are slightly better than DP-LoRA with 69.3%.
Since our benchmarking results from Section VI-C show that DP-LoRA achieves the best performance under DP-FL amongst all DP-PEFT methods, we run additionally experiments for LoRA under non-private federated learning to investigate the performance drop caused by providing DP guarantee. As we can see from Table IV and V, the average accuracy of 81.0% achieved by DP-LoRA is noticeably lower than the average accuracy of 88.2% under non-private FL.
@@
  • Overall, DP-LoRA outperforms other existing DP-PEFT methods under DP-FL for training large transformer-based models on-device. However, noticeable performance degradation can still be observed with a strong privacy budget of $\bm{\epsilon}$=2 and 1 million clients (over 7% in accuracy and over 17% in WER).
VI-D DP-DyLoRA
@@
Hyperparameter: We opt for the same federated learning and privacy parameters, and search over the same clipping thresholds and learning rates as in Section VI-C. As for parameters specific to DyLoRA, we set minimum and maximum ranks to $r_{min}$=1 and $r_{max}$=16. For DyLoRA, we perform evaluation at the server side at the end of every 10 rounds since each rank between $r_{min}$ and $r_{max}$ needs to be evaluated. On Sentiment140 and R8, we apply gradient clipping to DyLoRA under non-private FL as well since the model fails to converge otherwise. We additionally experiment with $r$={1, 8} with $\epsilon$=2 for both DP-LoRA and DP-DyLoRA for a more comprehensive comparison.
Results: As we can see from Table IV and V, DP-DyLoRA achieves an accuracy of 72.0% on Sentiment140 which is noticeably better than all existing DP-PEFT methods except for DP-LoRA with $r$=1 which achieves a marginally better accuracy of 72.5%. On the other datasets including R8, CIFAR-10, WikiArt, Speech Commands V1 and MINDS-14, DP-DyLoRA outperforms all other DP-PEFT methods including DP-LoRA with $r$={1, 8, 16}. Overall, DP-DyLoRA achieves a much better performance under DP-FL with an average accuracy of 86.4% as opposed to the 77.4%, 82.2% and 81.0% average accuracy achieved by DP-LoRA with $r$={1, 8, 16}, respectively. Similarly, for automatic speech recognition (ASR) on MINDS-14, the WERs of 81.6%, 53.1% and 51.7% achieved by DP-LoRA with $r$={1, 8, 16} respectively are significantly outperformed by DP-DyLoRA with a WER of 58.0%. We highlight the best accuracy or WER under DP-FL for each task and the best average accuracy across all five tasks in Table IV and V.
@@
Another interesting finding from our results shown in Table IV and V is that DyLoRA actually performs slightly worse than LoRA in non-private FL. We notice that DyLoRA training tends to be unstable in FL setting with a reasonably large learning rate. This is especially obvious during the early training stage. These results therefore show that the partial update of LoRA weights makes DyLoRA more sensitive to data heterogeneity. One possible remedy is to apply gradient clipping, which is mandatory under DP-FL. We show DyLoRA training under non-private FL on Sentiment140 both with and without gradient clipping applied in Appendix B.

### v4

@@ IV-G DyLoRA
$\begin{split}B_{b}&=B[:,1:b]\\ A_{b}&=A[1:b,:],\end{split}\vspace*{-4pt}$ (6)
@@
V DP-DyLoRA
In this section, we describe the proposed algorithm DP-DyLoRA for differentially private federated learning to train LoRA weights for a variable rank. Doing so both increases signal-to-noise ratio and saves privacy budget from hyperparameter searching. However, this runs up against a problem: the choice of rank $b$ would naturally be made per client, but this would not work with DP. If different clients were to send up different-length vectors, this would immediately break DP. If instead clients padded the statistics they sent up with zeros, this would decrease the signal-to-noise ratio on the highest ranks significantly.
@@ V DP-DyLoRA
DP-DyLoRA freezes all pre-trained weights and adds new trainable LoRA modules to make fine-tuning large pre-trained models more parameter-efficient. At client side, users train on their local data with a modified forward pass:
$h=W_{\text{pt}}x+\Delta Wx=W_{\text{pt}}x+B_{b}A_{b}x.$ (7)
Here, $W_{\text{pt}}$ is frozen and only $B_{b}$ and $A_{b}$ are trainable. The outputs of $W_{\text{pt}}x$ and $B_{b}A_{b}x$ are summed coordinate-wise and are given the same input for the forward pass. The updates to the trainable parameters $B_{b}$ and $A_{b}$ are then clipped and sent back to the server for aggregation and noise addition as in DP-FedAvg. The communication cost apart from the initial transfer of $W_{\text{pt}}$ is therefore equivalent to DP-LoRA with $r=\frac{r_{min}+r_{max}}{2}$ on average which is approximately half of that of DP-LoRA with $r=r_{max}$ assuming that $r_{min}=1$. Since the magnitude of the added noise grows with the number of parameters updated [14, 15], DP-DyLoRA also achieves higher signal-to-noise ratio than DP-LoRA under the same DP-FL setting. Meanwhile, the same level of model expressiveness is preserved as the model architecture and number of trainable parameters remain the same for the global model.
@@
V-A Security Analysis
Here, we analyse the security of our proposed method DP-DyLoRA and show that it satisfies $(\epsilon,\delta)$-DP.
At each communication round $t$, each sampled user in $C^{t}\subseteq\mathcal{K}$ sends model updates to the server after local training. Although we randomly sample the rank $b^{t}\in\{r_{min},r_{min}+1,\ldots,r_{max}\}$ to be trained at round $t$, each user sampled at the same communication round trains and shares parameters of the same rank $b^{t}$. Therefore, we use moments accountant [43] with Rényi Differential Privacy (RDP) [50] to privatise the shared model updates for a tight composition bound. For any $\alpha\in(1,\infty)$ and $\epsilon>0$, a randomised mechanism $\mathcal{M}$ satisfies $(\alpha,\epsilon^{\prime})$-RDP if for all neighbouring datasets $D$ and $D^{\prime}$, we have
$D_{\alpha}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))\triangleq\frac{1}{\alpha-1}\log\mathbb{E}\left(\frac{D(x)}{D^{\prime}(x)}\right)^{\alpha}\leq\epsilon^{\prime}.$ (8)
A randomised mechanism $\mathcal{M}$ that satisfies $(\alpha,\epsilon^{\prime})$-RDP also satisfies $(\epsilon,\delta)$-DP with
$\epsilon=\epsilon^{\prime}+\log(\frac{\alpha-1}{\alpha})-\frac{\log\alpha+\log\delta}{\alpha-1},$ (9)
for any $10<\delta<$ 1 [51].
Since we only consider training of large transformer-based models in this work, we use Gaussian mechanism [44] instead of Laplace mechanism [52] for achieving $(\epsilon,\delta)$-DP. This is because Gaussian mechanism allows the use of L2 sensitivity and the L2 sensitivity of a long vector is much smaller than its L1 sensitivity, leading to much less noise being added. To satisfy $(\epsilon,\delta)$-DP with Gaussian mechanism, noise of $\mathcal{N}(0,I\sigma^{2})$ is added to model updates with
$\vspace*{-2pt}\sigma=zS,\vspace*{-2pt}$ (10)
where $\sigma$ is used to compute Gaussian noise, $z$ is the noise multiplier calculated by the moments accountant for privacy parameters such as privacy budget and sampling rate and $S$ is the clipping threshold.
We further combine the Gaussian mechanism with privacy amplification via sampling [9, 10] to reduce the added Gaussian noise $\sigma$ to
$\sigma=\frac{zS}{|C|},$ (11)
where $|C|$ is the cohort size.
Tens of thousands of clients need to be sampled at each communication round for training large transformer-based models under DP-FL such as the ones we use for our experiments [53]. This number is reasonable in large-scale federated learning systems [19] but is difficult to achieve in simulation using current federated learning frameworks due to the amount of computation required [53]. Following [42, 53], we simulate the noise level of a larger cohort size $C_{large}$ with a smaller cohort size $C_{small}$ by
$\sigma^{+}=\frac{C_{small}}{C_{large}}\sigma^{\prime},$ (12)
where $\sigma^{+}$ is used to compute noise for larger cohort size and $\sigma^{\prime}$ is for smaller cohort size.
VI Experimental Setup
@@ VI-B Models
The same model is used for tasks of the same domain. Therefore, we use BERT-small [4] for experiments on Sentiment140 and R8, ViT-small [63] for CIFAR-10 and WikiArt, and DistilHuBERT [64, 65] for Speech Commands and MINDS-14. Despite the fact that these models are extremely small compared to state-of-the-art large language models with rapidly increasing sizes such as LLaMA [66, 67] with 70 billion parameters and GPT-4 [68] with 1.7 trillion parameters, models with over 20 million parameters are considered large for either on-device deployment or differentially private federated learning.
@@ VI-D Non-IID Partitioning
We therefore choose to utilise Dirichlet distribution to artificially achieve non-IID label distribution for the remaining four datasets, including R8, CIFAR-10, WikiArt and MINDS-14. For these datasets, we partition by drawing from a Dirichlet distribution with $\alpha=0.1$ by default following [69, 70, 71] which results in each client holding samples from very few classes.
VI-E Noise Mechanism
@@
VI-F Large Cohort Noise-Level Simulation
Following [42] and [53], we simulate the noise-level of larger cohort sizes with smaller ones. In practice, differentially private federated learning (DP-FL) is applied to systems with millions of clients [19]. However, it is infeasible to simulate this many clients due to resource constraints. Since a larger cohort size $C$ leads to less noise added for the same privacy guarantee, we simulate realistically large cohort size $C_{\text{large}}$ with smaller cohort size $C_{\text{small}}$. This makes our results more meaningful in terms of practical deployment of DP-FL. We follow the approach in [42] and [53] for achieving this. The noise $\sigma$ we use for simulation is then computed as $\sigma^{+}=\frac{C_{\text{small}}}{C_{\text{large}}}z_{\text{large}}\cdot S$ where $z_{\text{large}}$ is the noise multiplier calculated based on $C_{\text{large}}$.
VII Experiments
@@
VII-A Full Fine-tuning
DP-FL tends to degrade model performance due to the noise added to model updates. Previous works have proven that there is a proportional relationship between the number of updated parameters and the magnitude of the added noise [14, 15]. Hence, it is particularly challenging to train large models with differential privacy. Together with the potential non-independent and identically distributed (non-IID) data distribution challenge, large transformer models are likely to fail when learning under DP-FL.
@@ VII-B Data Heterogeneity
As shown in Figure 4 and 5, we train the model on each task with both IID and non-IID data partitioning using a single combination of client number and privacy budget taken from Q1. In Figure 4, we show results with $\alpha=[0.01,0.1,1000]$ for R8, CIFAR-10, WikiArt and MINDS-14 following [69, 70, 71] which are non-IID by drawing labels from a Dirichlet distribution. When $\alpha$ is set to $0.1$, there is hardly any difference from IID distribution with $\alpha=1000$ except for WikiArt on which the accuracy drops by approximately 3%. After further increasing the level of data heterogeneity by decreasing $\alpha$ to $0.01$, results are roughly the same on R8 and WikiArt. However, on CIFAR-10 and MINDS-14, model performance degrades by approximately 14% in accuracy and 10% in word error rate, respectively. This is caused by severe client drift due to data heterogeneity.
For Sentiment140 and Speech Commands which are both non-IID by natural factors, we can see from Figure 5 that the model achieves similar performance on the latter with both IID and non-IID distributions. This is likely due to the similar sample distributions by class. However, on the Sentiment140 dataset, the model performs noticeably better under IID data partitioning with approximately 4% improvement. Since Sentiment140 is a binary classification dataset, some clients may only hold samples of a single class even if it is not partitioned to have a non-IID label distribution. This leads to a relatively high level of data heterogeneity which is realistic in practice due to users having different interests and habits.
@@
VII-C Parameter-efficient Fine-tuning
Recent works [25, 16, 17] have started utilising parameter-efficient fine-tuning (PEFT) methods to fine-tune transformer models with differential privacy via both central and federate learning. Apart from the obvious benefits of lower computation and communication cost, the primary motivation originates from the proven fact that fewer trainable parameters lead to better privacy-utility trade-offs [14, 15].
We thereby start with comparing model training via full fine-tuning and PEFT with the same number of clients and privacy budget. Here, we use LoRA as an example of PEFT methods as it’s empirically proven to be superior to other popular PEFT methods on natural language understanding tasks in [25] and is used in [16] for research on DP-FL.
@@ VII-C Parameter-efficient Fine-tuning
Next, we benchmark exising PEFT methods under FL and DP-FL. We experiment with PEFT methods including Adapter, Compacter, BitFit and LoRA which are often considered in existing works on parameter-efficient fine-tuning [24, 25, 26]. We also include results for two popular variants of LoRA, namely LoHa [28] and AdaLoRA [29] to make our benchmark more comprehensive. Our benchmark covers three different domains including natural language understanding (NLU), computer vision (CV) and speech, similar to our previous experiments. Regarding datasets, we again choose to use Sentiment140/R8 for NLU, CIFAR-10/WikiArt for CV and Speech Commands V1/MINDS-14 for the speech domain.
Hyperparameter: For federated learning and privacy parameters, we use 1 million users, a subsampling rate of 1%, $\epsilon$=2, and $\delta$=1e-6 for all six datasets. For clipping threshold, we search over three values {0.1, 1.0, 10.0}. For Sentiment140, we train for 300 rounds and search over five learning rates {5e-2, 1e-1, 2e-1, 5e-1, 1e-0}. For R8, we train for 200 rounds and search over four learning rates {1e-1, 2e-1, 5e-1, 1e-0}. For CIFAR-10, we train for 100 rounds and search over four learning rates {2e-2, 5e-2, 1e-1, 2e-1}. For WikiArt, we train for 200 rounds and search over four learning rates {5e-2, 1e-1, 2e-1, 5e-1}. For Speech Commands V1, we train for 300 rounds and search over four learning rates {2e-1, 5e-1, 1e-0, 2e-0}. For MINDS-14, we train for 2000 rounds and search over four learning rates {2e-2, 5e-2, 1e-1, 2e-1}. Regarding parameters for DP-Adapter, DP-Compacter, DP-LoRA, DP-LoHa and DP-AdaLoRA, we choose to use $r$=16 for all methods and additionally $n$=8 for DP-Compacter and half of the initial rank as target rank for DP-AdaLoRA which are derived from previous works including [24, 25, 29].
@@
Results: Our benchmarking results covering DP-Adapter, DP-Compacter, DP-BitFit, DP-LoRA, DP-LoHa and DP-AdaLoRA across three different domains are shown in Table IV and V. On Sentiment140, DP-Adapter achieves the best accuracy of 70.7% which is marginally higher than 70.4% and 70.5% achieved by DP-LoRA and DP-AdaLoRA, respectively. On R8, DP-LoHa gives the best accuracy of 91.5%, followed by 90.0% for both DP-LoRA and DP-AdaLoRA. For image classification on CIFAR-10 and WikiArt, DP-AdaLoRA achieves the best accuracy of 95.0% for the former and DP-LoRA for the latter with an accuracy of 61.7%. On CIFAR-10, both DP-Adapter and DP-Compacter suffer from slow and unstable convergence which subsequently leads to much worse accuracy achieved. On Speech Commands V1, DP-LoRA once again achieves the best accuracy of 92.5%. The word error rates (WERs) of 58.9% and 59.7% achieved by DP-AdaLoRA and DP-LoHa respectively are better than the rest.
@@
  • Overall, DP-LoRA and its variants outperform other existing DP-PEFT methods under DP-FL for training large transformer-based models on-device. However, noticeable performance degradation can still be observed for DP-LoRA with a strong privacy budget of $\bm{\epsilon}$=2 and 1 million clients (over 7% in accuracy and over 17% in WER).
VII-D DP-DyLoRA
@@
Hyperparameter: We opt for the same federated learning and privacy parameters, and search over the same clipping thresholds and learning rates as in Section VII-C. As for parameters specific to DyLoRA, we set minimum and maximum ranks to $r_{min}$=1 and $r_{max}$=16. For DyLoRA, we perform evaluation at the server side at the end of every 10 rounds since each rank between $r_{min}$ and $r_{max}$ needs to be evaluated. On Sentiment140 and R8, we apply gradient clipping to DyLoRA under non-private FL as well since the model fails to converge otherwise. We additionally experiment with $r$={1, 8} with $\epsilon$=2 for both DP-LoRA and DP-DyLoRA for a more comprehensive comparison.
Results: As we can see from Table IV and V, DP-DyLoRA achieves an accuracy of 72.0% on Sentiment140 which is noticeably better than all existing DP-PEFT methods except for DP-LoRA with $r$=1 which achieves a marginally better accuracy of 72.5%. On CIFAR-10, the accuracy of 94.4% achieved by DP-DyLoRA is marginally lower than those of DP-LoHa and DP-AdaLoRA. On the other datasets including R8, WikiArt, Speech Commands V1 and MINDS-14, DP-DyLoRA outperforms all other DP-PEFT methods including DP-LoRA with $r$={1, 8, 16}. Overall, DP-DyLoRA achieves a much better performance under DP-FL with an average accuracy of 86.4% as opposed to the 77.4%, 82.2% and 81.0% average accuracy achieved by DP-LoRA with $r$={1, 8, 16}, respectively. Similarly, for automatic speech recognition (ASR) on MINDS-14, the WERs of 81.6%, 53.1% and 51.7% achieved by DP-LoRA with $r$={1, 8, 16} respectively are significantly outperformed by DP-DyLoRA with a WER of 58.0%. We highlight the best accuracy or WER under DP-FL for each task and the best average accuracy across all five tasks in Table IV and V.
@@
Another interesting finding from our results shown in Table IV and V is that DyLoRA actually performs slightly worse than LoRA in non-private FL. We notice that DyLoRA training tends to be unstable in FL setting with a reasonably large learning rate. This is especially obvious during the early training stage. These results therefore show that the partial update of LoRA weights makes DyLoRA more sensitive to data heterogeneity. One possible remedy is to apply gradient clipping, which is mandatory under DP-FL.